XYZ Company produces three products, A, B, and C. XYZ's plant capacity
is limited to 48,000 machine hours per year. The following information
is available for planning purposes:

                                    Product A       Product B       Product C
demand for next year .............  21,000 units   28,000 units   35,000 units
selling price per unit ...........    $80             $120            $160
direct material cost per unit ....    $24             $ 22            $ 36
direct labor cost per unit .......    $18             $ 50            $ 62
variable overhead cost per unit ..    $22             $ 25            $ 33 

It takes 0.80 machine hours to produce one unit of Product A; 1.20 machine
hours to produce one unit of Product B; and 1.60 machine hours to produce
one unit of Product C.

Calculate the total number of units of Product B that XYZ Company should
produce in order to maximize its net income.
Answer #1
Product A Product B Product C
Selling price per unit 80 120 160
Less: Variable costs
Direct material cost per unit 24 22 36
Direct labor cost per unit 18 50 62
Variable overhead cost per unit 22 25 33
Total Variable costs per unit 64 97 131
Contribution margin per unit 16 23 29
Divide by Machine hours per unit 0.80 1.20 1.60
Contribution margin per machine hour 20.00 19.17 18.13
Ranking in terms of production 1 2 3
Product A should be produced first
Machine hours for Product A 16800 =21000*0.80
Remaining Machine hours 31200 =48000-16800
Product B should be produced second
Units of Product B 26000 =31200/1.20
Total number of units of Product B = 26000